OpenStudy (anonymous):

if h(x)=x^2-5 find h(-3)

Nnesha (nnesha):

replace x by -3

OpenStudy (anonymous):

if I needed to find h(-2), this is what I would do: h(x)=x^2-5 h(-2)=(-2)^2-5 h(-2)=4-5 h(-2)=-1

Nnesha (nnesha):

 now replace that